solve using addition/elimination
x+y+z=0
x-y+4z=17
2x+y+z=-4

Answer by mananth(16946)

1 x + 1 y + 1 z = 0 -------------- 1
1 x + -1 y 4 z = 17 -------------- 2
2 x + 1 y + 1 z = -4 -------------- 3

consider equation 1 &2 Eliminate y
Add the two
2 x + 0 y + 5 z = 17 ------------- 4
consider equation 2 & 3 Eliminate y

Add the two
3 x + 0 y + 5 z = 13 -------------5 5
Consider (4) & (5) Eliminate x
Multiply 4 by -3
Multiply (5) by 2
we get
-6 x + -15 z = -51
6 x + 10 z = 26
Add the two
0 x + -5 z = -25
/ -5
z = 5
3 x + 5 z= 13
3 x + 25 = 13
x = -4
Plug x & z in eq (1)
1 x + 1 y + 1 z = 0 -------------- 1
-4 + 1 y + 5 = 0
1 y = -1
y = -1
x= -4 y= -1 z= 5